### v1

# Stronger Language Models Produce More Human-Like Errors

Abstract
Do language models converge toward human-like reasoning patterns as they improve? We provide surprising evidence that while overall reasoning capabilities increase with model sophistication, the nature of errors increasingly mirrors predictable human reasoning fallacies: a previously unobserved inverse scaling phenomenon. To investigate this question, we apply the Erotetic Theory of Reasoning (ETR), a formal cognitive framework with empirical support for predicting human reasoning outcomes. Using the open-source package PyETR, we generate logical reasoning problems where humans predictably err, evaluating responses from 38 language models across 383 reasoning tasks. Our analysis indicates that as models advance in general capability (as measured by Chatbot Arena scores), the proportion of their incorrect answers that align with ETR-predicted human fallacies tends to increase ($\rho=0.360,p=0.0265$). Notably, as we observe no correlation between model sophistication and logical correctness on these tasks, this shift in error patterns toward human-likeness occurs independently of error rate. These findings challenge the prevailing view that scaling language models naturally obtains normative rationality, suggesting instead a convergence toward human-like cognition inclusive of our characteristic biases and limitations, as we further confirm by demonstrating order-effects in language model reasoning.
1 Introduction
As language models rapidly advance in capabilities, they increasingly demonstrate sophisticated reasoning abilities across diverse domains [1, 2]. A naive prediction would have future language models achieving perfect rationality simply through scale. However, as these models become more capable of mimicking human language and problem-solving, do they also inherit our systematic reasoning errors?
Human reasoning exhibits robust patterns of fallacious inference that have been extensively documented in cognitive science [3, 4, 5, 6, 7]. These patterns are not random mistakes, but systematic deviations from normative standards that persist even among educated participants and that follow predictable patterns across diverse reasoning tasks. The study of these fallacies has been central to our understanding of human cognition and has profound implications for fields ranging from economics to medical decision-making.
Our work investigates whether increasingly sophisticated language models show a similar pattern of fallacious reasoning to humans. This question has both theoretical and practical significance. Theoretically, it provides insight into whether human-like reasoning errors emerge naturally from systems trained on human-generated text. Practically, it raises important considerations for AI systems deployed in contexts requiring reliable reasoning, such as medical diagnosis, legal analysis, or high-stakes decision support [8, 9].
To rigorously investigate this question, we leverage the Erotetic Theory of Reasoning (ETR) [10, 11, 12], a formal cognitive theory that precisely predicts human reasoning patterns across multiple domains. ETR posits that human reasoning operates by maintaining disjunctive alternatives and filtering these alternatives when taking on new information, a process that can systematically lead to error. ETR provides formal specifications of both how and when humans will make specific reasoning errors, allowing us to generate arbitrary reasoning problems with predictable failure patterns.
ETR has been implemented as PyETR [13], an open-source Python library. Using PyETR, we mechanically generate reasoning problems predicted to lead human-like reasoners into making errors. For our experiments we fix a diverse corpus of 383 such problems for which ETR predicts a specific fallacious response.
Our experimental investigation evaluates 38 language models ranging from smaller models (e.g., Mistral 7B Instruct v0.1) to state-of-the-art systems (e.g., GPT-4.5, Claude 3.7), using Chatbot Arena Elo as a proxy for model sophistication [14]. Our central finding is striking: as model capability increases, models become significantly more likely to commit human-like reasoning errors ($\rho=0.360,p=0.0265$).
This finding represents a notable example of an inverse scaling law in AI development—where model improvement on general metrics corresponds to decreased performance on specific desirable properties. While language models have demonstrated improved performance with scale on many reasoning tasks [1, 15], our work suggests that they may simultaneously become more susceptible to human-like reasoning fallacies. Beyond practical implications, our findings provide novel evidence relevant to the debate on whether LLMs develop internal reasoning processes analogous to human cognition, or simply mimic surface-level patterns from their training data.
Contributions
(1) We empirically demonstrate a significant inverse scaling relationship between language model sophistication and susceptibility to human-like reasoning fallacies. (2) We provide a theoretical analysis connecting our findings to broader questions about the emergence of human-like cognitive properties in language models. (3) With our methodology, we give the first substantial application of PyETR, an open-source computational implementation of the Erotetic Theory of Reasoning, to generate reasoning problems with predictable human-like failure patterns.
The remainder of this paper is organized as follows: Section 2 introduces the Erotetic Theory of Reasoning and our PyETR implementation. Section 3 details our experimental methodology. Section 4 presents our results. Section 5 discusses and concludes.
Related Work
Substantial prior literature in psychology and cognitive science documents irrationalities in human judgement and decision-making [3, 4, 5, 16]. In particular, the psychology of deductive inference reveals human propositional reasoning to be vulnerable to logically-irrelevant linguistic effects, like co-references between premises [6]. Mental model theory provides an explanatory paradigm for these results [17], though prior work on the Erotetic Theory of Reasoning matches and has replicated mental model predictions on deductive inference tasks [18, 19].
In the language model literature, prior studies find that while modern LLMs largely succeed at syllogistic inference [20, 21, 22], they exhibit human-like reasoning failures, including distractability [23], content effects [24], and order effects [21, 25]. Closest to our work, the first application of the Erotetic Theory to LLMs found the predictive utility of the theory to increase for larger model sizes [26], though this study was limited to OpenAI’s GPT model family [27, 28, 29] and did not make use of PyETR. Our implementation yields an endlessly regenerative and content-agnostic test of the Erotetic Theory for LLMs in a manner more resilient to data contamination.
2 The Erotetic Theory of Reason
Modern understanding of human reasoning reveals a fascinating duality: remarkable achievements coexist with persistent, predictable errors. The Erotetic Theory of Reasoning (ETR) provides a unified framework explaining both capacities through a single mechanism: humans construe problems as questions requiring answers, mentally representing disjunctive alternatives that they filter based on how well they match new information. This filtering process, where we prematurely dispose of alternatives before incorporating new data, enables cognitive efficiency but introduces systematic biases. Importantly, ETR demonstrates that reasoning success or failure depends not on separate cognitive systems but on whether one achieves “erotetic equilibrium,” a state where conclusions remain stable against potential follow-up questions.
At its core, ETR (full formalism in Appendix A) proposes a computational model where reasoning operates through three key mechanisms: (1) maintaining disjunctive alternatives representing possible answers to implicit questions, (2) filtering these alternatives by best-match with new information, potentially eliminating relevant possibilities prematurely, and (3) correcting potential errors by raising explicit questions that reintroduce necessary alternatives. This conceptually parsimonious framework has demonstrated remarkable empirical success in predicting systematic reasoning fallacies across propositional and first-order logic, probability theory, and decision-making. Mathematical proofs further establish that as reasoners approach complete erotetic equilibrium by raising appropriate questions, their conclusions converge to normative standards including logical soundness and completeness, Bayesian updating, and expected utility maximization, making ETR a compelling explanation for human reasoning as a unified cognitive capacity whose errors arise from computational constraints rather than irrationality.
Let us demonstrate the basics of ETR with the following reasoning problem, Example 49 in [10]. Premise 1 Either there is an ace in Mary’s hand and some other player has a king, or else there is a queen in John’s hand and some other player has a jack. Premise 2 Sally has a king. Question Does it follow that Mary has an ace? While the correct logical answer to the question is “not necessarily”, the conclusion “Mary has an ace” was endorsed by 60% of respondents (unpublished data following the methods of [11]).
To model the problem with logic, we would first choose a representation of the basic predicates, in this case $\mathsf{Ace}(x)$ for the predicate that (the card) $x$ is an ace, and $\mathsf{Has}(y,z)$ for the predicate that (player) $y$ holds (a card) $z$. The same applies when using ETR, where instead of logical formulas we have the central concept of view (see Appendix A) for the full mathematical formalism). It will suffice for now to use the (still precise) shorthand notation to express the premises and conclusion as ETR views: [^1] An additional aspect not present in ordinary logic is issue structure. This is notated with an asterisk on some occurrences of terms, indicative of being ‘at issue’ for the containing atomic proposition. The issue structure has no logical content, but plays a role in guiding ETR inferences. Determination of issue structure from cues in natural language or context is outside the scope of ETR, but a simple heuristic in this example is that the repetition of ‘Ace’ and ‘King’ suggests those concepts are at issue. P1 $\{\mathsf{Ace}(a^{*})\mathsf{Has}(\mathsf{Mary},a)\mathsf{King}(b^{*})\mathsf{Has}(c,b),\mathsf{Queen}(d)\mathsf{Has}(\mathsf{John},d)\mathsf{Jack}(e)\mathsf{Has}(f,e)\}$ P2 $\{\mathsf{King}(g^{*})\mathsf{Has}(\mathsf{Sally},g)\}$ Q $\{\mathsf{Ace}(h^{*})\mathsf{Has}(\mathsf{Mary},h)\}$
Each view is given as a disjunctive set of alternatives, separated by a comma. Each alternative is to be considered as a conjunctive set of atomic propositions, where for convenience we omit the set-brackets and commas and simply juxtapose elements. Many readers will recognize these as disjunctive normal forms, but in ETR there is no inherent ordering to the atoms in a conjunction or alternatives in the disjunction. The individual lowercase letters can be understood are as existentially quantified variables, with the scope of the quantification encompassing the entire disjunction (‘prenex form’). A view may also carry a supposition, written as a superscript (see Table 1 for examples).
PyETR, an open-source software package that serves as a calculator for ETR, can parse views from a similar text format.
from pyetr import View
[…]
Note the requirement to explicitly declare existentially quantified names (the letter E may be used in place of $\exists$). ETR posits some default procedures for ordinary reasoning, built out of some basic operations. The workhorse among the basic operations is Update, which update a current view with an incoming one. We can check in PyETR that ETR does predict a fallacy here, because Update discards the seemingly less relevant alternative where John has a queen.
$>$ v = p1.update(p2); print(v)
[…]
$>$ v.query(cc)
[…]
The actual default procedure of ETR involves a few more steps which become relevant in more complex problems. The full version is implemented in PyETR.
$>$ from pyetr.inference import default_procedure_does_it_follow
$>$ default_procedure_does_it_follow([p1,p2],cc)
True
3 Methodology
In our experiment, a reasoning problem is given by a list of views (the premises), and the problem is to answer the question “what if anything follows?”. We are most interested in reasoning problems where the ETR-predicted answer to this question is a logical fallacy.
3.1 Data processing
Table 1: Original bank of reasoning problems.
Problem Symbolic representation Natural language example
Modus ponens $\{R(x)\}^{\{Q(x)\}}$ If it rains, it’s wet.
$\{Q(x)\}$ It rains.
$\therefore\{R(x)\}$ Therefore, it’s wet.
Modus tollens $\{R(x)\}^{\{Q(x)\}}$ If the switch is on, then the lamp is lit.
$\{\overline{R(x)}\}$ The lamp is not lit.
$\therefore\{\overline{Q(x)}\}$ Therefore, the switch is not on.
Quantified modus ponens $\forall x\{R(x)\}^{\{Q(x)\}}$ All mammals have lungs.
$\forall x\{Q(x)\}^{\{P(x)\}}$ All dogs are mammals.
$\therefore\forall x\{R(x)\}^{\{P(x)\}}$ Therefore, all dogs have lungs.
Disjunction fallacy $\{Q(x)R(x),S(x)T(x)\}$ She is quiet and clever, or tall and athletic.
$\{Q(x)\}$ She is quiet.
$\therefore\{R(x)\}$ Therefore, she is clever.
We source our original bank of reasoning problems from examples presented in the Reason and Inquiry text [10]. These include basic “template” reasoning problems like modus ponens, modus tollens, and the disjunction fallacy as originally presented in Walsh and Johnson-Laird [6]. Table 1 presents the complete list of original problems. Not all reasoning problems take the form of valid syllogistic inferences; that is, not all conclusions from Erotetic Theory inference deductively follow from the premises.
3.2 Generative pipeline
We define a collection of mutation functions that give slight modifications to ETR view objects. These functions serve to expand our bank of reasoning problems by modifying existing problems by single premises. Table 2 presents the complete list of possible mutation rules.
Table 2: Logical Form Mutations
Mutation Type Description
Predicate Addition Introduces a new predicate symbol to the language.
Constant Addition Introduces a new constant symbol to the formal system.
Variable Addition Adds a new variable to the set of arbitrary objects.
Constant-to-Variable Substitution Replaces constants with $\forall$ or $\exists$ quantified variables.
Conjunctive Atom Insertion Conjoins new atomic formulas to existing states.
Disjunctive State Addition Disjunctively creates new states with atoms.
Atom Negation Applies or removes negation from atomic formulas.
Generating a new problem proceeds by iterative building a list of views (the premises in the problem). A new view is selected at random from the original problem bank and subjected to a random number of random mutation rules, it is added to the list if ETR predicts the problem admits a non-trivial answer. New views are added until all of the following stopping conditions are met. (1) The problem is the right size, meaning that summing the number of atoms in each view gives between 4 and 11. If 11 is exceeded, backtracking is used. (2) The ETR-predicted conclusion contains a single alternative (no disjunctions). (3) The ETR-predicted conclusion is a logical fallacy. To mitigate concerns about dataset-specific effects, we intentionally generated problems across diverse domains and structures. The PyETR-based generation ensures our problems test reasoning capacities rather than memorization of training examples.
3.3 Mapping to natural language
To evaluate language model reasoning in a more naturalistic context, we developed a systematic approach to convert formal logical statements into natural language prompts. We created 12 themes for natural language framings of logical problems, such as a researcher figuring out the properties of a novel element or a newly discovered creature. For each theme, we established consistent mappings between logical elements and thematic equivalents. The thematic conversion process serves multiple purposes: it ensures our results are robust against content effects [24]; it enables us to test whether the reasoning patterns are consistent in diverse contexts; and it mitigates potential data contamination, a significant problem for LLM evaluation [30, 31, 32, 33] by reformulating problems in novel scenarios unlikely to appear in training data.
Consider the following example of a disjunction fallacy problem in its formal representation: $\mathbf{P1}:\{Q(x)R(y),S(x)T(y)\},\mathbf{P2}:\{Q(x)\}$. In the “alchemy” theme, logical predicates (e.g., $A(x)$, $B(y)$) were mapped to thematic attributes (e.g., “is transmuting”, “is time-bending”), while logical variables (e.g., $x$, $y$) were mapped to imaginary thematic entities (e.g., “cosmic dust”, “vital mercury”). Using our alchemy theme, this was transformed into:
Im an alchemist studying mysterious substances in my laboratory. I need to understand their properties through logical reasoning. Heres what Ive discovered:
Cosmic dust is transmuting and vital mercury is time-bending, or cosmic dust is immortality-granting and vital mercury is spirit-affecting.
Also, we know that cosmic dust is immortality-granting.
What if anything follows? I do not have an intended answer in mind, and it is possible that nothing follows. Please be succinct and precise.
This exact prompt was sent to the LLM. It includes a thematic preamble and a natural language prompt to reason in an open-ended fashion with: “What if anything follows?” Each problem maintained the same prompt structure: an introduction establishing the thematic context, followed by the premises expressed in natural language, and concluding with a standardized question asking what follows from the premises. The prompt explicitly stated that nothing might follow, to avoid biasing models toward making unwarranted inferences [25].
3.4 Data Collection
Model Selection and Testing Conditions
We evaluated 38 language models of varying sophistication, including several specifically designed as reasoning models. Selection informally prioritised quantity, popularity of usage, and representation of a diverse range of scores. All models were accessed via OpenRouter and tested using LLMharness. Chatbot Arena scores were obtained from https://huggingface.co/spaces/lmarena-ai/chatbot-arena-leaderboard/resolve/main/elo_results_20250423.pkl, released under the Apache 2.0 license. To ensure a fair comparison across architectural differences, all reasoning models were allocated 2,400 thinking tokens during evaluation. This standardization was justified given our use of the Chatbot Arena Elo score as our primary capability metric, which already accounts for performance differences regardless of underlying mechanisms [14].
Response Processing and Filtering
Models provided conclusions in natural language, which we converted to PyETR format for evaluation. To ensure fair assessment of reasoning (not formatting), GPT-4.1-mini served as our translation layer without access to the original premises. Spot-checking confirmed translation fidelity.
Each model was evaluated on 383 reasoning tasks generated by PyETR. During analysis, we encountered occasional parse errors when processing model responses. To maintain data quality while preserving statistical power, models with parse error rates exceeding 20% across the test set were excluded from analysis, which resulted in the exclusion of one model (google__gemini-2.5-pro-preview-03-25). For the remaining models, individual responses that produced parse errors were excluded from analysis on a case-by-case basis rather than discarding all responses from that model.
Testing Framework
The full set of logical problems were framed as evals, using Eleuther’s Language Model Evaluation Harness [34], which handles evaling and metric collection against a wide range of models. OpenRouter [35] was used for all models, in order to have a unified API. This consumed less than \$1,000 of compute resources. Each model was given a token limit of 3,000 output tokens.
Key Measures
We collected the following measures to evaluate both the logical correctness of model responses and their alignment with human-like reasoning patterns, enabling our analysis of how these patterns correlate with model sophistication. A model’s answer was considered ‘logically correct’ if it was a logical consequence of the premises. This was tested using PySMT [36] to check whether the negation of the conclusion was inconsistent with the premises. LLM responses were classified as ‘ETR-predicted’ if ETR predicts an endorsement of that conclusion according to PyETR’s implementation of default_procedure_does_it_follow.
Statistical Tests
We selected Pearson correlation to capture linear relationships between variables while also reporting Spearman’s rank correlation to account for potentially non-linear monotonic relationships without assuming normality in the distribution of fallacy rates. For our premise-order intervention analysis, we employed two-proportion z-tests to rigorously evaluate whether reversing premise order produced statistically significant changes in fallacy rates. The z-test was specifically chosen because it allows us to determine if the observed differences in proportions (original vs. reversed premise order) represent genuine effects rather than random variation, providing a reliable measure of the significance of order effects across different model capabilities. This statistical approach enables us to quantify the extent to which models exhibit the same premise-order sensitivity documented in human reasoning studies.
4 Experimental Results
Our analysis reveals several notable patterns in how language models exhibit human-like reasoning fallacies. We report the results of Pearson correlation tests, as well as Spearman tests for the detection of non-parametric monotonic relationships. We consider $p\leq 0.05$ a suitable boundary for statistical significance.
Proportion of Errors that are Fallacies
We operationalize a “human-like fallacy” (or simply “fallacy”) as a response that is both ETR-predicted and logically incorrect. This definition captures instances where models make the same systematic reasoning errors that ETR predicts humans would make. Formally, for a given model $m$ and reasoning problem $p$, we define:
[…]
Our metric of interest is the fallacy rate, defined as the proportion of human-like fallacies to total logically incorrect answers for each model:
$\text{FallacyRate}(m)=\frac{\sum_{p\in P}\text{HumanLikeFallacy}(m,p)}{\sum_{p\in P}\neg\text{LogicallyCorrect}(m,p)}$
where $P$ is the set of all reasoning problems in our evaluation set. This measure represents the degree to which a model’s errors align with predictable human reasoning patterns rather than other forms of logical error. We find as model strength increases, so does the proportion of logically incorrect answers that are fallacies (Figure 1).
Figure 1: Over 38 models, a direct linear fit produces a significant moderate positive correlation ($r=0.364,p=\mathbf{0.0247}$), an exponential fit (log $y$-axis) displays a more confident Pearson correlation ($r=0.407,p=\mathbf{0.0113}$), as displayed above. As Spearman is ordinal, we have the same significant monotonic relationship in both cases ($\rho=0.360,p=\mathbf{0.0265}$).
Overall Correctness and other measures
Surprisingly, model strength as measured by LLM Elo score does not appear to have any correlation $(r=0.004,p=0.981)$, $(\rho=-0.04,p=0.777)$ with the model’s ability to produce correct answers on our dataset: see Table 3. Furthermore, exploratory analysis did not reveal any other statistically significant evidence of model Elo correlating or enjoying a monotonic ordinal relationship with absolute proportion of fallacies produced, whether exactly predicted by ETR or taking logical equivalences.
Table 3: Percentage of model responses which are logically correct
Mean $\sigma$ min $\mathcal{Q}(0.25)$ $\mathcal{Q}(0.5)$ $\mathcal{Q}(0.75)$ max
40.6% 16.7% 18.6% 31.0% 37.5% 44.8% 91.7%
Intervention study: Reversing Premise Order
While in classical logic the order in which premises are presented has no effect on normatively correct logical conclusions, an order-effect is commonly observed in human reasoners [37], where presenting premises in different orders elicits logically different responses. In addition to being a proxy measure for the non-classicality of human reasoners, this non-commutativity effect often blocks fallacies. We observe strong evidence that LLMs are similar in this regard: presenting the same questions with reversed premise order significantly blocks fallacy production. Table 4 reports a significant fallacy-blocking effect from reversing premises on a majority of models.
Table 4: 38 models selected for testing, listed by Chatbot Arena Score (general). The right-most column reports the percentage of fallacies per-model blocked (answer becomes logically correct) by reversing premise order, along with the results of a two-proportion $z$-test for each model to gauge the significance of the blocking effect.
Provider Model Elo Fallacies blocked by reversal
OpenAI chatgpt-4o-latest 1408 46.15%, (z = 3.24, p = 1.18e-03)
OpenAI gpt-4.5-preview 1398 68.00%, (z = 3.03, p = 2.44e-03)
Google gemini-2.5-flash-preview 1393 37.27%, (z = 3.52, p = 4.27e-04)
DeepSeek deepseek-chat-v3-0324 1373 38.78%, (z = 2.28, p = 2.27e-02)
Google gemma-3-27b-it 1342 20.39%, (z = 1.77, p = 7.60e-02)
OpenAI o1-mini 1304 50.00%, (z = 0.58, p = 5.63e-01)
Anthropic claude-3.7-sonnet 1292 50.00%, (z = 4.23, p = 2.35e-05)
xAI grok-2-1212 1288 22.68%, (z = 1.91, p = 5.58e-02)
Anthropic claude-3.5-sonnet 1283 65.08%, (z = 4.74, p = 2.09e-06)
OpenAI gpt-4o-mini-2024-07-18 1272 36.36%, (z = 2.00, p = 4.59e-02)
Google gemini-flash-1.5 1271 29.69%, (z = 1.98, p = 4.78e-02)
Mistral mistral-large-2407 1251 37.60%, (z = 3.85, p = 1.18e-04)
Meta llama-3.1-70b-instruct 1248 24.32%, (z = 1.74, p = 8.14e-02)
Anthropic claude-3-opus 1247 32.63%, (z = 2.79, p = 5.20e-03)
Mistral mistral-small-24b-instruct-2501 1217 34.53%, (z = 3.81, p = 1.40e-04)
Microsoft phi-4 1205 44.68%, (z = 3.91, p = 9.34e-05)
Anthropic claude-3-sonnet 1201 53.49%, (z = 3.04, p = 2.40e-03)
Google gemma-2-9b-it 1192 55.17%, (z = 2.55, p = 1.07e-02)
Cohere command-r-plus-04-2024 1190 37.25%, (z = 2.22, p = 2.64e-02)
OpenAI gpt-4-0314 1186 41.67%, (z = 2.07, p = 3.81e-02)
Anthropic claude-3-haiku 1179 56.67%, (z = 3.91, p = 9.20e-05)
OpenAI gpt-4 1163 54.05%, (z = 2.85, p = 4.43e-03)
Meta llama-3-8b-instruct 1152 23.81%, (z = 1.55, p = 1.21e-01)
Mistral mistral-medium 1148 43.90%, (z = 2.38, p = 1.75e-02)
Mistral mixtral-8x22b-instruct 1148 42.86%, (z = 2.12, p = 3.43e-02)
Anthropic claude-2.0 1132 32.81%, (z = 3.40, p = 6.65e-04)
Anthropic claude-2.1 1118 54.79%, (z = 4.28, p = 1.91e-05)
Mistral mixtral-8x7b-instruct 1114 37.50%, (z = 1.49, p = 1.36e-01)
OpenAI gpt-3.5-turbo-0125 1106 39.42%, (z = 3.61, p = 3.06e-04)
Meta llama-3.2-3b-instruct 1103 20.00%, (z = 0.77, p = 4.41e-01)
Nous Research nous-hermes-2-mixtral-8x7b-dpo 1084 66.67%, (z = 2.91, p = 3.63e-03)
DeepSeek deepseek-chat 1077 35.00%, (z = 2.29, p = 2.19e-02)
Mistral mistral-7b-instruct-v0.2 1072 38.46%, (z = 1.11, p = 2.66e-01)
OpenAI gpt-3.5-turbo-1106 1068 88.46%, (z = 4.36, p = 1.33e-05)
Meta llama-3.2-1b-instruct 1054 32.43%, (z = 1.60, p = 1.10e-01)
Microsoft phi-3-mini-128k-instruct 1037 38.16%, (z = 2.88, p = 3.96e-03)
AllenAI olmo-7b-instruct 1015 45.45%, (z = 1.23, p = 2.18e-01)
Mistral mistral-7b-instruct-v0.1 1008 66.67%, (z = 1.42, p = 1.55e-01)
5 Conclusion and Discussion
Interpretation of Key Findings Our results reveal a statistically significant trend: as language models advance in capability (measured by Chatbot Arena Elo), their logical errors increasingly align with human errors predicted by ETR. Critically, while stronger language models demonstrate improved capabilities across most benchmarks, we observe a systematic increase in human-like fallibility: an inverse scaling trend that persists across model families, architectures, and training paradigms.
Moreover, we demonstrate order-effects for LLM reasoning analogous to those observed in humans. This finding suggests that beyond merely emulating modal human reasoning patterns, LLMs exhibit signs of mirroring human cognition at deeper levels, including characteristic fallacies.
Methodological Considerations and Limitations While traditional ablation studies might seem appropriate for investigating this phenomenon, they face significant limitations in the current landscape of language models: the diversity of modern architectures, training approaches such as distillation  [38, 39, 40], and reasoning capabilities with test-time scaling  [41, 42, 43] means that simple parameters like model size may no longer serve as reliable proxies for model capability. Our use of Chatbot Arena scores provides a performance-based metric that naturally incorporates a diverse range of models, reflecting real-world ability rather than architectural specifics, similar to Ruan et al. [44].
Though the correlation strength ($\rho=0.360$) may appear moderate, it is robust across diverse model architectures and training paradigms, suggesting a fundamental relationship rather than artifact. The statistical significance ($p=0.0265$) exceeds conventional thresholds despite our conservative analytic approach.
While our results demonstrate correlation between model capability and human-like errors, we acknowledge limitations in attributing causality. Alternative explanations include the possibility that stronger models are increasingly trained on human-produced reasoning traces that themselves contain these fallacies. The absence of improved logical accuracy with model capability may reflect ceiling effects in our test set or fundamental limitations in current training paradigms.
Theoretical Implications These findings challenge the assumption that scaling alone leads to more rational systems. Instead, we observe convergence toward human-like reasoning patterns including characteristic fallacies, suggesting that current language models may be developing internal processes that functionally mirror human cognitive shortcuts rather than idealized rational agents. If models increasingly adopt human-like reasoning patterns, including our characteristic errors, this suggests the emergence of cognitive structures functionally similar to human reasoning, rather than alternative solution methods that happen to produce similar outputs.
Connection to Inverse Scaling Laws and Practical Applications Our findings represent an interesting case of an inverse scaling law [45], where improvement on general metrics correlates with increasing susceptibility to specific reasoning fallacies. This phenomenon has both theoretical significance and practical applications for AI development.
Our results provide a foundation for several promising research directions. For benchmark development, our approach offers a systematic method for generating reasoning problems where human-like fallacies are predicted, creating evaluation sets that could complement existing reasoning benchmarks. For capabilities, the identified fallacy patterns could serve as targeted training examples to fortify models against human-like reasoning errors while maintaining their general capabilities. As a tool for assessing engineering choices, our methodology provides a quantitative framework for evaluating the effectiveness of various interventions aimed at improving model reasoning, from prompt engineering to architectural modifications.
Broader Implications This inverse scaling phenomenon has profound implications for AI alignment. As models become more capable, anticipating and mitigating human-like reasoning errors becomes increasingly important, especially in high-stakes contexts requiring reliable reasoning, such as medical diagnosis, legal analysis, or decision support [8, 9].
We see these results as opening new avenues for understanding both artificial and human reasoning. As language models continue to advance, this work provides a foundation for developing systems that combine human-like understanding with more robust reasoning capabilities: systems that can successfully reason as we do while avoiding the pitfalls that characterize human cognition.
Acknowledgments and Disclosure of Funding
The authors wish to thank the Cosmos Institute for their continued support.
Appendix A A formal presentation of the Erotetic Theory
ETR is formulated in [10] in a set-theoretic language, and we reproduce the full definitions here, with the explicit permission of the authors. We remark that Definitions 1 and 11 contain the parameters of the theory, i.e. the basic atomic propositions, the constants and function symbols, and a set logical and extralogical axioms. The central definition is that of a View, Definition 9. Following that are definitions of the basic operations of ETR, in particular Update (Definition 23) and Query (Definition 30) whose use in PyETR were demonstrated in Section 2.
Definition 1 (Basic objects).
Let $\mathbf{A},\mathbf{F},\{?\}$, $\mathbf{P}^{\top}$, $\mathbf{P}^{\bot}$ be pairwise disjoint countable sets.
  • ($\mathbf{F}$ arity)
    $\alpha:\mathbf{F}\to\mathbb{N}$. Write $f\in\mathbf{F}$ as $f^{k}$ when $\alpha(f)=k$.
  • (Constants)
    $\exists w_{\in\mathbf{F}}\ \alpha(w)=0$.
  • ($\mathbf{P}$ polarity)
    Let $N:\mathbf{P}^{\top}\to\mathbf{P}^{\bot}$ be a bijection. Write $\bar{P}$ for $N(P_{\in\mathbf{P}^{\top}})$.
    Let $\mathbf{P}=\mathbf{P}^{\top}\cup\mathbf{P}^{\bot}$.
  • ($\mathbf{P}$ arity)
    $\alpha^{\prime}:\mathbf{P}\to\mathbb{N}$. Write $P$ as $P^{k}$ for $\alpha^{\prime}(P)=k$. $\alpha^{\prime}(P)=\alpha^{\prime}(\bar{P})$.
  • (Identity)
    ${=^{2}}\in\mathbf{P}$.
Definition 2 (Terms $\mathbf{T}$).
$\mathbf{A}\subseteq\mathbf{T}$. If $f^{0}\in\mathbf{F}$, then $\langle f^{0}\rangle\in\mathbf{T}$. If $f^{k}\in\mathbf{F}\wedge\{t_{1}\ldots t_{k}\}\subseteq\mathbf{T}$, then $\langle f^{k},\langle t_{1}\ldots t_{k}\rangle\rangle\in\mathbf{T}$.
Definition 3 (Atoms $\mathcal{A}$).
If $P^{k}\in\mathbf{P}$ and $\vec{\tau}=\langle t_{1}\ldots t_{k}\rangle$ for $t_{i}\in\mathbf{T}$, then $\langle P^{k},\vec{\tau}\rangle\in\mathcal{A}$ (abbrv. $P\vec{\tau}$). $|\vec{\tau}|=\{t_{1}\ldots t_{k}\}$.
Definition 4 (States $\mathbb{S}$).
[…]
Definition 5 (Dependency relations $\mathcal{R}$).
Let $\Gamma\in\mathcal{P}(\mathbb{S})$. Let $U,E\subseteq\mathbf{A}(\Gamma)$ and $D\subseteq E\times U=\{\langle e,u\rangle:e\in E\wedge u\in U\}$. Define $\langle U,E,D\rangle\in\mathcal{R}_{\Gamma}$ iff
  • (Bipartite)
    $U\cap E=\emptyset\wedge D\subseteq E\times U$, and
  • (Matryoshka)
    For all $u,u^{\prime}\in U$$\{e\in E:\langle e,u\rangle\in D\}\subseteq\{e\in E:\langle e,u^{\prime}\rangle\in D\}$ or $\{e\in E:\langle e,u^{\prime}\rangle\in D\}\subseteq\{e\in E:\langle e,u\rangle\in D\}$.
[…]
Definition 6 (Open terms $\mathbf{T_{1}}\textbf{and atoms }\mathcal{A}_{1}$).
?$\in\mathbf{T}_{1}$ and $\langle f^{k},\langle t_{1},\ldots,t_{k}\rangle\rangle\in\mathbf{T}_{1}$ whenever $f^{k}\in\mathbf{F}$ and for some $1\leq i\leq k$, $t_{i}\in\mathbf{T}_{1}$ and $t_{j}\in\mathbf{T}$ for $j\neq i$. $\langle P^{k},\langle t_{1},\ldots,t_{k}\rangle\rangle\in\mathcal{A}_{1}$, whenever $P^{k}\in\mathbf{P}$ and for some $1\leq i\leq k$, $t_{j}\in\mathbf{T}$ for all $j\neq i$ and $t_{i}\in\mathbf{T}_{1}$.
Definition 7 (Issue structures $\mathbb{I}$).
$I\in\mathbb{I}_{\Gamma}$ iff $I$ consists of pairs $\langle t,x\rangle$, s.t. $x\in\mathcal{A}_{1}$ and $x[t/?]\in\mathcal{A}(\Gamma)$.
Definition 8 (Issue matches $M_{IJ}$).
For $I,J\in\mathbb{I}_{\Gamma}$, define,
$M_{IJ}=\{\langle t_{1},t_{2}\rangle:\exists x(\langle t_{1},x\rangle\in I\wedge(\langle t_{2},x\rangle\in J\vee\langle t_{2},\bar{x}\rangle\in J))\}$
where, for $x=\langle P^{k},\langle t_{1},\ldots,t_{k}\rangle\rangle$, $\bar{x}=\langle\bar{P}^{k},\langle t_{1},\ldots,t_{k}\rangle\rangle$.
Definition 9 (Views $\mathbb{V}$).
For $\Gamma,\Theta$ finite subsets of $\mathbb{S}$, $R\in\mathcal{R}_{\Gamma\cup\Theta}$, $I\in\mathbb{I}_{\Gamma\cup\Theta}$,
$\langle\Gamma,\Theta,R,I\rangle\in\mathbb{V}$ (abbreviated $\Gamma^{\Theta}_{RI}\in\mathbb{V}$).
Write $\top$ for $\{0\}^{\{0\}}_{0_{\mathcal{R}}\emptyset}$ and $\bot$ for $\emptyset^{\{0\}}_{0_{\mathcal{R}}\emptyset}$.
Definition 10 (Commitments $\mathbb{C}$).
For $C\subseteq\mathbb{V},G\in\mathbb{V}$, $\langle C,G\rangle\in\mathbb{C}$.
Definition 11 (Primitive absurd states $\mathbb{K}$).
Let $\mathbb{K}_{\subseteq\mathbb{S}}$ contain, $\forall t,t^{\prime}_{\in\mathbf{T}}\forall p_{\in\mathcal{A}}\forall x_{\in\mathcal{A}_{1}}$, at least:
  • (LNC)
    $\{p,\bar{p}\}$,
  • (Aristotle)
    $\{{\neq}tt\}$,
  • (Leibniz)
    $\{{=}tt^{\prime},x[t/?],\bar{x}[t^{\prime}/?]\}$.
Definition 12 ($\mathcal{R}$ Restriction).
$[R]_{X}=\langle U_{R}\cap X,E_{R}\cap X,D_{R}\cap((E_{R}\cap X)\times(U_{R}\cap X))\rangle$.
$[R]_{\Gamma}=[R]_{\mathbf{A}(\Gamma)}$.
Given $\langle\Gamma,\Theta,R,I\rangle$, we allow ourselves to write $\langle\Gamma,\Theta,[R],I\rangle$ for $\langle\Gamma,\Theta,[R]_{\Gamma\cup\Theta},I\rangle$.
Definition 13 ($\mathbb{I}$ Restriction).
Within a quadruple $\langle\Gamma,\Theta,R,[I]\rangle$, let
$[I]=\{\langle t,x\rangle:\langle t,x\rangle\in I\wedge x[t/?]\in\mathcal{A}(\Gamma\cup\Theta)\}$
Definition 14 ($\mathcal{R}$ Algebra).
Let $R\rtimes S=\langle U_{R}\cup U_{S},E_{R}\cup E_{S},D_{R}\cup D_{S}\cup E_{S}\times U_{R}\rangle$ and let
$0_{\mathcal{R}}\bowtie 0_{\mathcal{R}}=0_{\mathcal{R}}$.
Let
$R\bowtie S=\langle U_{0},E_{0},\emptyset\rangle\rtimes([R]_{\mathbf{A}(R)-(E_{0}\cup U_{0})}\bowtie[S]_{\mathbf{A}(S)-(E_{0}\cup U_{0})})$
where $E_{0}=\{e_{\in E_{R}\cup E_{S}}:\forall u.\langle e,u\rangle\notin D_{R}\cup D_{S}\}$ and
$U_{0}=\{u_{\in U_{R}\cup U_{S}}:\forall e\notin E_{0}.\langle e,u\rangle\notin(E_{R}\times U_{R}-D_{R})\cup(E_{S}\times U_{S}-D_{S})\}$.
Definition 15 (Product).
$\Gamma^{\Theta}\otimes\Delta^{\Psi}=\big{(}\{\gamma_{\in\Gamma}\cup\delta_{\in\Delta}:\exists\psi_{\in\Psi}(\psi\subseteq\gamma)\}\cup\{\gamma_{\in\Gamma}:\neg\exists\psi_{\in\Psi}(\psi\subseteq\gamma)\}\big{)}^{\Theta}$
$(\Gamma\otimes\Delta)^{\{0\}}=\Gamma^{\{0\}}\otimes\Delta^{\{0\}}$
$\bigotimes_{i\in\mathbf{P}}\Delta_{i}^{\Psi_{i}}=\{0\}^{\{0\}}\otimes\Delta_{1}^{\Psi_{1}}\otimes\ldots\otimes\Delta_{n}^{\Psi_{n}}$
$\Gamma^{\Theta}_{RI}\otimes^{T}\Delta^{\Psi}_{SJ}=(\Gamma^{\Theta}\otimes\Delta^{\Psi})_{[(T\bowtie R)\bowtie(T\bowtie S)][I\cup J]}$
$\Gamma^{\Theta}_{RI}\otimes\Delta^{\Psi}_{SJ}=\Gamma^{\Theta}_{RI}\otimes^{0_{\mathcal{R}}}\Delta^{\Psi}_{SJ}$
$\bigotimes^{T}_{i\in\mathbf{P}}{\Delta_{i}}^{\Psi_{i}}_{S_{i}J_{i}}=\top\otimes^{T}{\Delta_{1}}^{\Psi_{1}}_{S_{1}J_{1}}\otimes^{T}\ldots\otimes^{T}{\Delta_{n}}^{\Psi_{n}}_{S_{n}J_{n}}$
Definition 16 (Sum).
$\Gamma^{\Theta}_{RI}\oplus^{T}\Delta^{\Theta}_{SJ}=(\Gamma\cup\Delta)^{\Theta}_{[(T\bowtie R)\bowtie(T\bowtie S)][I\cup J]}$
$\Gamma^{\Theta}_{RI}\oplus\Delta^{\Theta}_{SJ}=\Gamma^{\Theta}_{RI}\oplus^{0_{\mathcal{R}}}\Delta^{\Theta}_{SJ}$
$\bigoplus^{T}_{i\in\mathbf{P}}{\Delta}^{\Psi_{i}}_{S_{i}J_{i}}=\emptyset^{\Psi}_{0_{\mathcal{R}}\emptyset}\oplus^{T}{\Delta_{1}}^{\Psi}_{S_{1}J_{1}}\oplus^{T}\ldots\oplus^{T}{\Delta_{n}}^{\Psi}_{S_{n}J_{n}}$
Definition 17 ($\mathcal{A}$ Answer potential).
$\Gamma[\Delta]^{\mathcal{A}P}=|\mathcal{A}(\Gamma)\cap\mathcal{A}(\Delta)|$
Definition 18 (Answer).
$\Gamma^{\Theta}_{RI}[\Delta^{\{0\}}_{SJ}]^{A}=(\operatorname{argmax}\limits_{\gamma\in\Gamma}\ \Delta[\{\{p\}:p\in\gamma\}]^{\mathcal{A}P})^{\Theta}_{[R][I]}$
Definition 19 (Negation).
$[\Gamma^{\Theta}_{RI}]^{N}=(\Theta^{\{0\}}\otimes([\Gamma]^{N})^{\{0\}})_{[R]^{N}[I]^{N}}$
$[\Gamma]^{N}=\bigotimes_{\gamma\in\Gamma}\hat{\gamma}^{\{0\}}$
$\hat{\gamma}=\big{\{}\{\bar{p}\}:p_{\in\gamma}\big{\}}$
$\bar{p}=\bar{F}\vec{\tau}\text{ if }p=F\vec{\tau};$ $\bar{p}=F\vec{\tau}\text{ if }p=\bar{F}\vec{\tau}$
$[R]^{N}=\langle E_{R},U_{R},\{\langle a,b\rangle\in U_{R}\times E_{R}:\langle b,a\rangle\notin D_{R}\}\rangle$
$[I]^{N}=I\cup\{\langle t,\bar{x}\rangle:\langle t,x\rangle\in I\}$
Definition 20 (Novelty).
$a\in\mathbf{A}$ is novel during an inference step $\langle C,G\rangle[D]^{O}$ if $a\notin\mathbf{A}(G)\cup\mathbf{A}(D)$ and $a$ has not yet appeared in the computation of the conclusion of the step. A function $\nu:X\to\mathbf{A}$, where $X\subseteq\mathbf{A}$ is finite, is novel during an inference step if each $\nu(x)$ is novel and $\nu$ is injective, and then $[\nu]_{X}$ denotes the simultaneous substitution $[\nu(a_{1})/a_{1},\ldots,\nu(a_{k})/a_{k}]$, where $X=\{a_{1},\ldots,a_{k}\}$. Let $(-)^{\mathrm{nov}(X)}$ stand for $(-)[\nu]_{X}$ where $\nu:X\to\mathbf{A}$ is novel.
Definition 21 (Substitution).
Let $Z(T,a)=\{u\in U_{T}:u\triangleleft_{T}a\}\cup\{e\in E_{T}:e\lesssim_{T}a\}-\{a\}=\{a_{1},\ldots,a_{k}\}$. For $\mathbf{A}(\Gamma)\cap\mathbf{A}(\Theta)=\emptyset$,
$\mathrm{Sub}^{T}_{\langle t,a\rangle}(\Gamma^{\Theta}_{I})=\langle\Gamma[\nu_{1}]_{Z(T,a)}[t/a],\Theta,[T\rtimes([T]_{Z(T,a)}[\nu_{1}]_{Z(T,a)})],I[\nu_{1}]_{Z(T,a)}[t/a]\rangle,$
where $\nu_{1}:Z(T,a)\to\mathbf{A}$ is novel.
Definition 22 (Merge).
For either $\mathbf{A}(R)\cap\mathbf{A}(S)=\emptyset$ or $[R]_{\Delta\cup\Psi}=S$,
$\Gamma^{\Theta}_{RI}[\Delta^{\Psi}_{SJ}]^{M}=\bigoplus^{R\bowtie S}_{\gamma\in\Gamma}\Big{(}\{\gamma\}^{\Theta}_{RI}\otimes^{R\bowtie S}\Delta^{\Psi}_{SJ}\otimes^{R\bowtie S}\bigotimes^{R\bowtie S}_{\langle t,u\rangle\in M^{\prime}_{IJ}(\gamma)}\mathrm{Sub}^{R\bowtie S}_{\langle t,u\rangle}(\Delta^{\{0\}}_{J})\Big{)}$,
where $M^{\prime}_{IJ}(\gamma)=\{\langle t,u\rangle\in M_{IJ}:u\in U_{S}\wedge\exists\psi_{\in\Psi}(\psi[t/u]\subseteq\gamma\wedge\psi\not\subseteq\gamma)\}$
Definition 23 (Update).
For $D\in C$, but with all arbitrary objects novelised,
$\langle C,\Gamma^{\Theta}_{RI}\rangle[D]^{\circlearrowright}=\langle C,\Gamma^{\Theta}_{RI}[D]^{U}[D]^{E}[D]^{A}[D]^{M}\rangle$
Definition 24 (Universal product).
For $\mathbf{A}(\Gamma)\cap\mathbf{A}(\Theta)=\emptyset$ and either $\mathbf{A}(R)\cap\mathbf{A}(S)=\emptyset$ or $[R]_{\Delta}=S$,
$\Gamma^{\Theta}_{RI}[\Delta^{\{0\}}_{SJ}]^{U}=\{0\}^{\Theta}_{RI}\otimes^{R\bowtie S}\bigotimes^{R\bowtie S}_{\langle u,t\rangle\in M^{\prime}_{IJ}}\mathrm{Sub}^{R\bowtie S}_{\langle t,u\rangle}(\Gamma^{\{0\}}_{I})$, where
$M^{\prime}_{IJ}=\{\langle u,t\rangle:\langle u,t\rangle\in M_{IJ}\wedge u\in U_{R}-\mathbf{A}(\Theta)\}\neq\emptyset$
Definition 25 (Reorient).
For $\Delta^{\Psi}_{SJ}\in(C\cup\{\Gamma^{\Theta}_{RI}\})$ and $J^{\prime}\in\mathbb{I}_{\Delta\cup\Psi}$,
$\langle C,\Gamma^{\Theta}_{RI}\rangle[\Delta^{\Psi}_{SJ^{\prime}}]^{R}=\langle C,\Delta^{\Psi}_{SJ^{\prime}}\rangle$
Definition 26 (Existential sum).
Let the following conditions be met:
  • (1)
    $\mathbf{A}(\Gamma)\cap\mathbf{A}(\Theta)=\emptyset$ and either $\mathbf{A}(R)\cap\mathbf{A}(S)=\emptyset$ or $[R]_{\Delta}=S$.
  • (2)
    $M^{\prime}_{IJ}=\{\langle e,t\rangle:\langle e,t\rangle\in M_{IJ}\wedge e\in E_{R}-\mathbf{A}(\Theta\cup\Delta)\wedge\neg\exists x(\langle e,x\rangle\in D_{R})\}\neq\emptyset$.
Then,
$\Gamma^{\Theta}_{RI}[\Delta^{\{0\}}_{SJ}]^{E}=\Gamma^{\Theta}_{RI}\oplus^{R\bowtie S}\bigoplus^{R\bowtie S}_{\langle e,t\rangle\in M^{\prime}_{IJ}}\\ \mathrm{Sub}^{R\bowtie S}_{\langle t,e\rangle}\big{(}(\bigcup_{\gamma\in\{\gamma_{\in\Gamma}:e\in\mathbf{A}(\gamma)\}}(\{\gamma\}\cup\big{\{}\{x_{\in\gamma}:e\notin\mathbf{A}(x)\}\cup\delta:\\ \delta\in\bigotimes_{x\in B(\gamma,I,e)}\{\{x\},\{\bar{x}\}\}\wedge\delta\not\subseteq\gamma\big{\}}))^{\Theta}_{I}\big{)},$
where for $e\in E_{R}$ and $\gamma\in\Gamma$, let $B(\gamma,I,e)=\{x[e/?]\in\gamma:\langle e,x\rangle\in I\}$.
Definition 27 (Division).
If $\forall\delta_{\in\Delta}\exists\psi_{\in\Psi}\exists\gamma_{\in\Gamma}(\delta\subseteq\gamma\wedge\psi\subseteq\gamma)$, then
$\Gamma^{\Theta}_{RI}\oslash\Delta^{\Psi}_{SJ}=\{\gamma\oslash_{\Gamma}\Delta^{\Psi}:\gamma\in\Gamma\}^{\Theta}_{[R][I]}$
$\gamma\oslash_{\Gamma}\Delta^{\Psi}=\gamma-\iota\delta(\delta\in\Delta\wedge\delta\subseteq\gamma\wedge\exists\psi_{\in\Psi}(\psi\subseteq\gamma))$
Definition 28 (Factor).
For $\Delta^{\Psi}_{SJ}\neq\bot$,
$\Gamma^{\Theta}_{RI}[\Delta^{\Psi}_{SJ}]^{F}=\{\gamma[\Delta^{\Psi}]^{F}:\gamma\in\Gamma\}^{\Theta}_{[R][I]}$ where
$\gamma[\Delta^{\Psi}]^{F}=(\gamma\oslash_{\Gamma}\Delta^{\Psi})\cap\bigcap\{\gamma\oslash_{\Gamma}(\Delta^{\Psi}[t/a]):\langle t,a\rangle\in M_{IJ}\wedge a\in U_{S}\}$, and where we let $\gamma^{\prime}\cap\bigcap\emptyset=\gamma^{\prime}$.
For $\Delta^{\Psi}_{SJ}=\bot$,
$\Gamma^{\Theta}_{RI}[\bot]^{F}=\big{\{}\gamma\in\Gamma:\neg\exists\delta\in\mathbb{K}.\delta\subseteq\gamma\big{\}}^{\Theta}_{[R][I]}$.
Definition 29 (Matryoshka level ordering).
For existentials, we have $e\lesssim_{R}e^{\prime}$ iff $\forall\langle e^{\prime},u\rangle\in D_{R}(\langle e,u\rangle\in D_{R})$. For both existentials and universals, let the relation $\triangleleft_{R}$ be the transitive closure of the relation made up by the ordered pairs in $D_{R}\cup(E_{R}\times U_{R}-D_{R})^{\mathrm{op}}$, where $op$ reverses the order of pairs in a set of pairs.
Definition 30 (Query).
For $U_{S}\subseteq U_{R}$ and $M^{\prime}_{IJ}=\{\langle t,e\rangle\in M_{IJ}:e\in E_{S}\backslash E_{R}\}$, we define
$\Gamma^{\Theta}_{RI}[\Delta^{\Psi}_{SJ}]^{Q}=$
($(\{0:\neg\exists\delta_{\in\Delta}\exists\gamma.\Phi(\gamma,\delta)\}\cup\{\delta\in\Delta:\exists\gamma_{\in\Gamma}.\Phi(\gamma,\delta)\})^{\Theta}_{[R\bowtie\langle U_{R},E_{S}\backslash E_{R},D_{S^{\prime}}\rangle][I\cup J]}$
where $\Phi(\gamma,\delta)\leftrightarrow\exists\psi_{\in\Psi}\exists n_{\geq 0}\exists\langle t_{1},e_{1}\rangle,\ldots,\langle t_{n},e_{n}\rangle_{\in M^{\prime}_{IJ}}\forall i,j$
$(\psi\cup\delta[t_{1}/e_{1},\ldots,t_{n}/e_{n}]\subseteq\gamma\wedge e_{i}=e_{j}\to i=j)$;
and $D_{S^{\prime}}=D_{1}\cup D_{2}\cup D_{3}\cup D_{4}\cup D_{5}\cup D_{6}$ is constructed by taking: original dependencies from the internal argument
$D_{1}=[D_{S}]_{\mathbf{A}/E_{R}},$
extra dependencies for multiple terms substituted with same $e$
$D_{2}=\{\langle e_{m},u\rangle:\exists m\exists m^{\prime}\in M^{\prime}_{IJ}(e_{m}=e_{m^{\prime}}\wedge t_{m}\neq t_{m^{\prime}})\wedge u\in U_{R}\},$
dependencies resulting from complex terms being substituted
$D_{3}=\{\langle e_{m},u\rangle:\langle t_{m},e_{m}\rangle\in M^{\prime}_{IJ}\wedge u\in U_{R}(t_{m})\},$
$D_{4}=\{\langle e_{m},u\rangle:\langle t_{m},e_{m}\rangle\in M^{\prime}_{IJ}\wedge e\in E_{R}(t_{m})\wedge\langle e,u\rangle\in R\},$
$D_{5}=\{\langle e_{m},u\rangle:\langle t_{m},e_{m}\rangle\in M^{\prime}_{IJ}\wedge u\in U_{R}\wedge\forall u^{\prime}_{\in U_{R}(D_{3}\cup D_{4})}(u^{\prime}\triangleleft_{R}u)\},$
additional dependencies necessary to preserve the dependency order of existentials in $S$
$D_{6}=\{\langle e,u\rangle:e,e^{\prime}\in E_{S}-E_{R}\wedge e\lesssim_{S}e^{\prime}\wedge\\ (\forall m,m^{\prime}\in M^{\prime}_{IJ}(e_{m}=e^{\prime}\wedge e_{m^{\prime}}=e^{\prime})\to t_{m}=t_{m^{\prime}})\wedge\langle e^{\prime},u\rangle\in D_{1}\cup D_{2}\cup D_{3}\cup D_{4}\cup D_{5}\}.$
Definition 31 (Wh-Query).
Provided that $U_{S}\subseteq U_{R}$,
$\Gamma^{\Theta}_{RI}[\Delta^{\Psi}_{SJ}]^{W}=\Big{(}\big{\{}0:\exists\gamma\in\Gamma.\neg\exists\delta_{\in\Delta}.\Phi(\gamma,\delta)\big{\}}\cup\big{\{}\delta\in\Delta:\exists\gamma_{\in\Gamma}.\Psi(\gamma,\xi)\big{\}}\Big{)}^{\Theta}_{[R][I]},$ where $\Psi(\gamma,\xi)$ iff
$\exists\psi_{\in\Psi}\exists n_{\geq 0}\exists\langle t_{1},e_{1}\rangle,\ldots,\langle t_{n},e_{n}\rangle_{\in M^{\prime}_{IJ}}\exists\delta_{\in\Delta}.\\ \big{(}\xi\cup\psi\subseteq\gamma\wedge\xi=\delta[t_{1}/e_{1},\ldots,t_{n}/e_{n}]\\ \wedge\forall i,j.(e_{i}=e_{j}\to i=j)\big{)}.$
Definition 32 (Inquire).
(O) If $\mathbf{A}(\Gamma\cup\Theta)\cap\mathbf{A}(\Delta\cup\Psi)=\emptyset\wedge\mathbf{A}(\Delta)\cap\mathbf{A}(\Psi)=\emptyset$, then
$\Gamma^{\Theta}_{RI}[\Delta^{\Psi}_{SJ}]^{I}=\Gamma^{\Theta}_{RI}\otimes\Big{(}\Delta^{\Psi}_{SJ}\oplus\big{(}\{0\}^{\Psi}_{[S][J]}\otimes Nov([\Delta^{\{0\}}]^{N}_{[[S]^{N}][J]^{N}})\big{)}\Big{)}[\bot]^{F}$,
where $Nov()$ uniformly replaces all arbitrary objects in its scope by novel ones.
(I) If $\mathbf{A}(\Delta\cup\Psi)\subseteq\mathbf{A}(\Gamma\cup\Theta)$, then
$\Gamma^{\Theta}_{RI}[\Delta^{\Psi}_{SJ}]^{I}=\Gamma^{\Theta}_{RI}\otimes\big{(}\Delta^{\Psi}_{0_{\mathcal{R}}J}\oplus([\Delta]^{N}_{0_{\mathcal{R}}[J]^{N}})^{\Psi}\big{)}[\bot]^{F}$.
Definition 33 (Suppose).
(O) If $\mathbf{A}(\Gamma\cup\Theta)\cap\mathbf{A}(\Delta\cup\Psi)=\emptyset$, then
$\Gamma^{\Theta}_{RI}[\Delta^{\Psi}_{SJ}]^{S}=\Gamma_{[R\bowtie R^{\prime}][I\cup I^{\prime}]}^{\Theta^{\prime}}[\Delta^{\Psi}_{SJ}]^{U}[\Delta^{\Psi}_{SJ}]^{E}[\Delta^{\Psi}_{SJ}]^{A}[\Delta^{\Psi}_{SJ}]^{M}$, where
${\Theta^{\prime}}^{\{0\}}_{R^{\prime}I^{\prime}}=\Theta^{\{0\}}_{RI}\otimes Nov(\Delta^{\Psi}_{[S]^{N}J}[\ ]^{D})$,
where $Nov()$ uniformly replaces all arbitrary objects in its scope by novel ones.
(I) If $\mathbf{A}(\Delta)\subseteq\mathbf{A}(\Gamma\cup\Theta)$ and $[R]_{\Delta}=S$,
then $\Gamma^{\Theta}_{RI}[\Delta^{\{0\}}_{SJ}]^{S}=\Gamma^{\Theta\otimes\Delta}_{RI}[\Delta^{\{0\}}_{SJ}]^{U}[\Delta^{\{0\}}_{SJ}]^{E}[\Delta^{\{0\}}_{SJ}]^{A}[\Delta^{\{0\}}_{SJ}]^{M}$
Definition 34 (Depose).
$\Gamma^{\Theta}_{RI}[\top]^{D}=\langle\Gamma^{\{0\}}\oplus([\Theta]^{N})^{\{0\}},R,[[I]^{N}\cup I]\rangle$.
Definition 35 (Commit).
$\langle C,\Gamma^{\Theta}_{RI}\rangle[\top]^{C}=\langle C\cup\{\Gamma^{\Theta}_{RI}\},\Gamma^{\Theta}_{RI}\rangle$
Definition 36 (Inference).
An inference $\mathcal{I}$ is a finite sequence of inference steps defined as follows, relative to an inference state consisting of a commitment $\langle C,G\rangle\in\mathbb{C}$. If $E\in\mathbb{V}$ and
$O\in\{[D]^{\circlearrowright},[D]^{S},[D]^{Q},[D]^{F},[D]^{R},[D]^{I},[D]^{D},[D]^{E},[D]^{U},[D]^{C}\}$, then $O$ is an inference step applicable to $\langle C,G\rangle$, whose result is an inference state in $\mathbb{C}$. We write $C\vdash_{\mathrm{ETR}}{}E$ if and only if there is an erotetic theory inference $\mathcal{I}$ s.t. $\langle C,\top\rangle\mathcal{I}=\langle C^{\prime},E^{\prime}\rangle$, where $E^{\prime}$ is identical to $E$ up to uniform replacement of $\mathbf{A}$ objects.